OpenStudy (anonymous):

find the axis of symmetry and vertex of y=-3xsquared + 4

OpenStudy (anonymous):

vertex(0,-4)

OpenStudy (anonymous):

axis of symmetry is x = -b/2a a = -3 b = 0 x = 0/2(-3) so x = 0 Axis of symmetry is x = 0 To find the y, put in x = 0 into the original equation y = -3*0^2 + 4 y = 4 vertex (0,4)